OpenStudy (anonymous):

Solve the Quadratic Equation: 3x^2-24x=-36

OpenStudy (anonymous):

first divide everything by 3

OpenStudy (anonymous):

\[x^2-8x=-12\] \[(x-4)^2=-12+16=4\] \[x-4=2\] or \[x-4=-2\]\[x=6\] or \[x=2\]

OpenStudy (anonymous):

UHH WRONG ACTUALLY ^ its 6.6 and 1.7